OpenStudy (anonymous):

Jenna has two fish tanks in the shape of rectangular prisms. Each of the sides of the second fish tank is 5 times greater than the sides of the first fish tank. If Jenna fills both the fish tanks completely with water, the volume of water in the second fish tank will be 125 times the volume of water in the first fish tank 10 times the volume of water in the first fish tank 25 times the volume of water in the first fish tank 5 times the volume of water in the first fish tank

jimthompson5910 (jim_thompson5910):

each dimension of the second tank is 5 times greater so the second tank will a volume that's 5*5*5 = 5^3 = 125 times greater than the first one

jimthompson5910 (jim_thompson5910):

general rule if each dimension of a solid is k times greater than other solid, then its volume will be k^3 times greater